### v1

@@ 2.5 Charge conservation
The charge deposition scheme is placed after the particle push and before the field solver. In this way, it has full access to the initial and final positions of each macro-particle, i.e. ${r}^{n}$ and ${r}^{n+1}$, as well as their half-step velocities, ${u}^{n+1/2}$. The first allows us to deposit the initial and final charge density in the grid nodes, i.e. $\rho_{i,j}^{n}$ and $\rho_{i,j}^{n+1}$ are known, following Appendix B. Reconstructing the charged current density from the half-step velocities and the charge densities to satisfy equation (28) is the challenge we need to overcome. The solution to this problem was inspired in the seminal work by Villasenor & Buneman (1992) (hereafter VB) that suggested the computation of the charged current density directly from inverting equation (28), thus enforcing it by construction. As a particle moves from its initial position ${r}^{n}$ to its final position ${r}^{n+1}$, it may cross grid cell boundaries, hence continuously changing its macro-particle shape. To address this, we split the inter-cell motion into several intra-cell trajectories and tackle each split individually, as shown in Fig.2. Without loss of generality, we will describe how to retrieve the charged current density contribution from a single macro-particle given that its motion is bound to a single cell. The algorithm starts by splitting the divergence operator into its radial and meridional components:
(${J}\right)=\left(\@vec{\nabla}\cdot\@vec{J}\right)^{\hat{r}}+\left(\@vec{\nabla}\cdot\@vec{J}\right)^{\hat{\theta}}.$ (29)
The VB method suggests that each component of (29) can be used to evaluate $J^{\hat{r}}_{i+1/2,j}$ and $J^{\hat{\theta}}_{i,j+1/2}$ by computing the temporal derivatives of the charge density when the particle moves purely along the radial (meridional) direction with an average meridional (radial) position. Here, we take a different approach due to the continuous change in the particle shape. We assume that the particle shape changes more significantly along the radial direction; thus, (${J})^{\hat{\theta}}$ cannot be evaluated at a fixed average radius. To overcome this issue, we compute it via
(${J}\right)^{\hat{\theta}}=\left(\@vec{\nabla}\cdot\@vec{J}\right)-\left(\@vec{\nabla}\cdot\@vec{J}\right)^{\hat{r}},$ (30)
@@
(${J}\right)^{n+1/2}_{i,j}$ $=-\frac{\partial\rho_{i,j}}{\partial t}\Biggr{|}^{r^{n+1},\theta^{n+1}}_{r^{n},\theta^{n}}=-\frac{\rho_{i,j}\left(r^{n+1},\theta^{n+1}\right)-\rho_{i,j}\left(r^{n},\theta^{n}\right)}{\Delta t},$ (31)
(${J}\right)^{\hat{r},n+1/2}_{i,j}$ $=-\frac{\partial\rho_{i,j}}{\partial t}\Biggr{|}^{r^{n+1},\bar{\theta}}_{r^{n},\bar{\theta}}=-\frac{\rho_{i,j}\left(r^{n+1},\bar{\theta}\right)-\rho_{i,j}\left(r^{n},\bar{\theta}\right)}{\Delta t},$ (32)
@@
(${J}\right)^{\hat{r}}_{i,j}$ $=\frac{A^{\hat{r}}_{i+1/2,j}J^{\hat{r}}_{i+1/2,j}-A^{\hat{r}}_{i-1/2,j}J^{\hat{r}}_{i-1/2,j}}{V_{i,j}}=\frac{A^{\hat{r}}_{i+1/2,j}J^{\hat{r}}_{i+1/2,j}}{V_{i,j}},$ (33)
(${J}\right)^{\hat{\theta}}_{i,j}$ $=\frac{A^{\hat{\theta}}_{i,j+1/2}J^{\hat{\theta}}_{i,j+1/2}-A^{\hat{\theta}}_{i,j-1/2}J^{\hat{\theta}}_{i,j-1/2}}{V_{i,j}}=\frac{A^{\hat{\theta}}_{i,j+1/2}J^{\hat{\theta}}_{i,j+1/2}}{V_{i,j}},$ (34)
where $V_{i,j}$ is the volume of the cell evaluated at the cell node, see Appendix A. To justify the disappearance of the second terms in equations (33) and (34), we use the scheme presented in Fig. 2. Note that each split deposits the charged current density in the grid cell its trajectory is bound to. Finally, we obtain the poloidal components of the current by combining equations (32) with (33), and (30) with (34):
$J_{i+1/2,j}^{\hat{r},n+1/2}$ $=\frac{V_{i,j}}{A_{i+1/2,j}^{\hat{r}}}\left(\@vec{\nabla}\cdot\@vec{J}\right)^{\hat{r},n+1/2}_{i,j}$ (35)
$J_{i,j+1/2}^{\hat{\theta},n+1/2}$ $=\frac{V_{i,j}}{A_{i,j+1/2}^{\hat{\theta}}}\left(\@vec{\nabla}\cdot\@vec{J}\right)^{\hat{\theta},n+1/2}_{i,j}.$ (36)
@@
$J_{i,j}^{\hat{\phi},n+1/2}=\bar{\rho}_{i,j}\left(\bar{\alpha}\frac{u^{\hat{\phi},n+1/2}}{\Gamma\left(\@vec{u}^{n+1/2}\right)}-\bar{\beta}^{\hat{\phi}}\right),$ (37)
@@
The logical spherical coordinate system has an advantage over the standard spherical coordinate system as it admits a Cartesian-like grid. Nevertheless, this methodology has shown machine-precision charge conservation even for the standard spherical coordinate system with a non-uniform poloidal grid.
@@ 2.6 Boundary conditions
We now discuss the boundary conditions available in the OSIRIS-GR particle-in-cell framework. OSIRIS allows for two possible domains: the full or the half domain. These two setups differ in their meridional extension. The first extends from the north pole (at $\theta=0$) to the south pole (at $\theta=\pi$), while the second runs from the north pole to the equator (at $\theta=\pi/2$). The field and particle boundary conditions change accordingly.
@@ 2.6.1 Field boundary conditions
The radial boundaries also take two options: star or open. The star radial boundary condition mimics the surface of a perfect rotating spherical conductor threaded by a dipolar magnetic field. We impose on the interior guard cells and on the surface components (i.e. $E^{\hat{\phi}}$, $E^{\hat{\theta}}$ and $B^{\hat{r}}$) the interior electromagnetic solution of an isolated compact neutron star (equations (144)-(147) in Torres et al. (2023)). We compute the guard cell values of the current components such that their values are progressively filtered (i.e. with an increasing filter order starting from 0 for the first radial cell to the desired order $n$ for the $(n+1)^{th}$ radial cell).
The open radial boundary, placed at the outer frontier of the domain, is a Mur-like absorbing boundary condition (Mur 1981) designed to absorb outwardly propagating electromagnetic waves. This condition for an arbitrary field component $\Phi$ in flat spacetime reads:
@@
which is the Sommerfeld radiation condition in spherical coordinates (Novak & Bonazzola 2004; Espinoza et al. 2014). In Appendix C, we present the extension of equation (40) for curved spacetime and use it to correct the fields that lie in the exterior boundary (i.e. $E^{\hat{\phi}}$, $E^{\hat{\theta}}$, and $B^{\hat{r}}$). This open boundary condition was already used in Torres et al. (2023) and demonstrated efficient absorption of transient electromagnetic waves launched from a compact neutron star in vacuum. We take the same procedure for the charged current density as in the inner radial boundary.
@@ 2.7 Particle injection mechanism
OSIRIS allows for several plasma injection mechanisms. Here we limit the discussion to the ones used in the simulations presented in upcoming sections. We distinguish two types of simulations, with or without heuristic pair production (Cruz et al. 2022). For the latter, we inject pair plasma from the stellar surface at a constant rate with number density $n_{\mathrm{inj}}$ every timestep. Electrons and positrons are initialized in corotation and with a parallel velocity of $0.5\leavevmode\nobreak\ [\mathrm{c}]$. This injection procedure is limited by the local value of the magnetization parameter $\sigma_{*}$, given by:
@@
where $B_{*}$ is the local amplitude of the magnetic field, $\Gamma$ is the Lorentz factor of the injected particles, and $n_{\pm}$ is the number density of positrons or electrons. Setting a minimum value for $\sigma_{*}$ yields a maximum value of the injected plasma density. With the parallel component of the velocity, we can control the injected current provided to the neutron star’s magnetosphere. Then, the magnetosphere extracts the sign of the charges required to drive the magnetospheric currents and sustain the global torsion of the magnetic field lines. The goal is to mimic the plasma supply provided by the vacuum work function and the pair-producing cascade close to the stellar surface, i.e. the polar cap gap, while still populating both the closed field line region and the return current. Previous works using global particle-in-cell simulations have already taken this approach as it is a computationally efficient way to populate the magnetosphere without requiring expensive quantum electrodynamics (QED) processes (Cerutti et al. 2015; Philippov et al. 2015a). Nevertheless, these processes are already implemented in the OSIRIS framework, to model QED cascades in future ultra-intense laser experiments (Grismayer et al. 2016, 2017) or in local neutron star polar caps (Cruz et al. 2021).
The second approach is to sustain a cold corotating atmosphere close to the surface. We characterize the atmosphere by its value at the surface, $n_{0}$, and the scale height of the exponentially decaying profile, $\sigma_{atm}$:
@@
We ensure the non-depletion of this atmosphere up to $k_{atm}$ standard deviations from its base by injecting $n_{\mathrm{inj}}$ plasma number density every timestep when required. This approach differs from the previous one as it allows the magnetosphere to regulate the flux of plasma extracted from the corotating atmosphere. This extraction mechanism would then seed the cascading QED process. To reduce the computational cost of each simulation, we model the cascade with a heuristic pair production (Chen & Beloborodov 2014; Philippov et al. 2015a; Cruz et al. 2022, 2023). This model for the operation of the vacuum gap defines a threshold energy $\gamma_{thr}$ for a particle to pair produce. Particles that satisfy this condition can pair produce if their radial position corresponds to the range $R_{*}\leq r\leq R_{\mathrm{PP}}$. Also, this threshold energy has a polar dependence to inhibit pair production close to the polar axis, where the magnetic conversion does not occur due to the infinite magnetic field line curvature. The secondary pair is emitted at the same place and with the same momentum direction as the parent with a Lorentz factor of $\gamma_{pair}/2$. These parameters are user-defined and can be adjusted to study different magnetospheric solutions. This model has shown to be a good model for the polar cap and the outer gaps (Chen & Beloborodov 2014; Philippov et al. 2015a; Bransgrove et al. 2022; Hu & Beloborodov 2022; Cruz et al. 2022, 2023).
@@ 2.8 Simulation initialization and transient phase
Simulations start in vacuum with a prescribed general-relativistic dipolar magnetic field (Rezzolla et al. 2001; Torres et al. 2023) threading the neutron star surface. Equivalently, the angular velocity of the star can be either in full rotation or gradually spun to its desired value. For the first scenario, the general-relativistic electric field (Torres et al. 2023) must be prescribed at $t=0$. In the second scenario, the rotation imparted by the stellar boundary condition (see Sect. 2.6.1) generates the self-consistent electric field by launching Alfvénic torsion waves (Cerutti et al. 2015). At $t=0$, we initiate the plasma injection mechanism described in Sect. 2.7. Due to the high magnetization, the extracted charges from the surface/atmosphere are accelerated along field lines, generating a poloidal current that opens up the field lines that cross the light-cylinder radius, located at $R_{LC}\equiv c/\Omega_{*}$. Beyond this distance, particles can no longer co-rotate with the field lines anchored to the stellar surface as they would need to co-rotate at speeds faster than the speed of light. These particles then sustain the torsion of the field lines that acquire a toroidal component. In the aligned rotator case, when the magnetic and rotation axes are aligned, the extracted current at the poles is negative due to the predominant outflow of electrons. The magnetosphere self-organizes to close the electric current, which generates the return current. This positive electric current neutralizes the charge density of the neutron star and flows between the closed and open field lines in a Y-shaped structure. Beyond the Y-point, the current sheet sustains the toroidal and poloidal magnetic field reversal at the equator. The magnetospheric solution starts to form in the back of the outwardly propagating Alfvénic torsion wave. We have assumed that after two rotation periods of the star, the magnetospheric solution has converged to the quasi-steady state.

### v2

@@ 2.5 Current deposition scheme
The charge deposition scheme is placed after the particle push and before the field solver. In this way, it has full access to the initial and final positions of each macro-particle, i.e. $\boldsymbol{r}^{n}$ and $\boldsymbol{r}^{n+1}$, as well as their half-step velocities, $\boldsymbol{u}^{n+1/2}$. The first allows us to deposit the initial and final charge density in the grid nodes, i.e. $\rho_{i,j}^{n}$ and $\rho_{i,j}^{n+1}$ are known, following B. Reconstructing the charged current density from the half-step velocities and the charge densities to satisfy equation (28) is the challenge we need to overcome. The solution to this problem was inspired in the seminal work by Villasenor and Buneman (1992) (hereafter VB) that suggested the computation of the charged current density directly from inverting equation (28), thus enforcing it by construction. As a particle moves from its initial position $\boldsymbol{r}^{n}$ to its final position $\boldsymbol{r}^{n+1}$, it may cross grid cell boundaries, hence continuously changing its macro-particle shape. To address this, we split the inter-cell motion into several intra-cell trajectories and tackle each split individually, as shown in Fig.2. Without loss of generality, we will describe how to retrieve the charged current density contribution from a single macro-particle given that its motion is bound to a single cell. The algorithm starts by splitting the divergence operator into its radial and meridional components:
$\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)=\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{r}}+\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{\theta}}.$ (29)
The VB method suggests that each component of (29) can be used to evaluate $J^{\hat{r}}_{i+1/2,j}$ and $J^{\hat{\theta}}_{i,j+1/2}$ by computing the temporal derivatives of the charge density when the particle moves purely along the radial (meridional) direction with an average meridional (radial) position. Here, we take a different approach due to the continuous change in the particle shape. We assume that the particle shape changes more significantly along the radial direction; thus, $(\boldsymbol{\nabla}\cdot\boldsymbol{J})^{\hat{\theta}}$ cannot be evaluated at a fixed average radius. To overcome this issue, we compute it via
$\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{\theta}}=\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)-\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{r}},$ (30)
@@
$\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{n+1/2}_{i,j}$ $=-\frac{\partial\rho_{i,j}}{\partial t}\Biggr{|}^{r^{n+1},\theta^{n+1}}_{r^{n},\theta^{n}}=-\frac{\rho_{i,j}\left(r^{n+1},\theta^{n+1}\right)-\rho_{i,j}\left(r^{n},\theta^{n}\right)}{\Delta t},$ (31)
$\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{r},n+1/2}_{i,j}$ $=-\frac{\partial\rho_{i,j}}{\partial t}\Biggr{|}^{r^{n+1},\bar{\theta}}_{r^{n},\bar{\theta}}=-\frac{\rho_{i,j}\left(r^{n+1},\bar{\theta}\right)-\rho_{i,j}\left(r^{n},\bar{\theta}\right)}{\Delta t},$ (32)
@@
$\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{r}}_{i,j}$ $=\frac{A^{\hat{r}}_{i+1/2,j}J^{\hat{r}}_{i+1/2,j}-A^{\hat{r}}_{i-1/2,j}J^{\hat{r}}_{i-1/2,j}}{V_{i,j}}=\frac{A^{\hat{r}}_{i+1/2,j}J^{\hat{r}}_{i+1/2,j}}{V_{i,j}},$ (33)
$\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{\theta}}_{i,j}$ $=\frac{A^{\hat{\theta}}_{i,j+1/2}J^{\hat{\theta}}_{i,j+1/2}-A^{\hat{\theta}}_{i,j-1/2}J^{\hat{\theta}}_{i,j-1/2}}{V_{i,j}}=\frac{A^{\hat{\theta}}_{i,j+1/2}J^{\hat{\theta}}_{i,j+1/2}}{V_{i,j}},$ (34)
where $V_{i,j}$ is the volume of the cell evaluated at the cell node, see A. To justify the disappearance of the second terms in equations (33) and (34), we use the scheme presented in Fig. 2. Note that each split deposits the charged current density in the grid cell its trajectory is bound to. Finally, we obtain the poloidal components of the current by combining equations (32) with (33), and (30) with (34):
$J_{i+1/2,j}^{\hat{r},n+1/2}$ $=\frac{V_{i,j}}{A_{i+1/2,j}^{\hat{r}}}\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{r},n+1/2}_{i,j}$ (35)
$J_{i,j+1/2}^{\hat{\theta},n+1/2}$ $=\frac{V_{i,j}}{A_{i,j+1/2}^{\hat{\theta}}}\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{\theta},n+1/2}_{i,j}.$ (36)
@@
$J_{i,j}^{\hat{\phi},n+1/2}=\bar{\rho}_{i,j}\left(\bar{\alpha}\frac{u^{\hat{\phi},n+1/2}}{\Gamma\left(\boldsymbol{u}^{n+1/2}\right)}-\bar{\beta}^{\hat{\phi}}\right),$ (37)
@@
The current deposit algorithm presented above conserves charge to machine precision independently of the selected coordinates (polar spherical or logical spherical, see C). The main difference between both is, in fact, the choice of the cell center position, modifying the effective particle shape. The logical spherical coordinate system possesses a uniform Cartesian-like grid and asymmetric particle shape, thus solving the cell expansion issue characteristic of the symmetric particle shape and non-uniform grid case (i.e., polar spherical case). The latter case requires some particles to deposit charge and current densities in more than two consecutive cells (Cruz et al., 2023).
@@ 2.6 Boundary conditions
We now discuss the boundary conditions available in the OSIRIS-GR particle-in-cell framework. As the OSIRIS code allows for two possible domains (full or half meridional domain), the field and particle boundary conditions change accordingly.
@@ 2.6.1 Field boundary conditions
The radial boundaries also take two options: star or open. The star radial boundary condition mimics the surface of a perfect rotating spherical conductor threaded by a dipolar magnetic field. We impose on the interior guard cells and on the surface components (i.e. $E^{\hat{\phi}}$, $E^{\hat{\theta}}$ and $B^{\hat{r}}$) the interior electromagnetic solution of an isolated compact neutron star (equations (144)-(147) in Torres et al. (2023)). We compute the guard cell values of the current components such that their values are progressively filtered (i.e. with an increasing filter order starting from 0 for the first radial cell to the desired order $n$ for the $(n+1)^{th}$ radial cell - see D for more details).
The open radial boundary, placed at the outer frontier of the domain, is a Mur-like absorbing boundary condition (Mur, 1981) designed to absorb outwardly propagating electromagnetic waves. This condition for an arbitrary field component $\Phi$ in flat spacetime reads:
@@
which is the Sommerfeld radiation condition in spherical coordinates (Novak and Bonazzola, 2004; Espinoza et al., 2014). In E, we present the extension of equation (40) for curved spacetime and use it to correct the fields that lie in the exterior boundary (i.e. $E^{\hat{\phi}}$, $E^{\hat{\theta}}$, and $B^{\hat{r}}$). This open boundary condition was already used in Torres et al. (2023) and demonstrated efficient absorption of transient electromagnetic waves launched from a compact neutron star in vacuum. We take the same procedure for the charged current density as in the inner radial boundary.
@@ 2.7 Particle injection mechanism
OSIRIS allows for several plasma injection mechanisms. Here we limit the discussion to the ones used in the simulations presented in upcoming sections. We distinguish two types of simulations, with or without heuristic pair production (Cruz et al., 2022). For the latter, we inject pair plasma from the stellar surface at a constant rate with number density $n_{\mathrm{inj}}$ every timestep. Electrons and positrons are initialized in corotation and with a parallel velocity of $0.5~{}[\mathrm{c}]$. This injection procedure is limited by the local value of the magnetization parameter $\sigma_{*}$, given by:
@@
where $B_{*}$ is the local amplitude of the magnetic field, $\Gamma$ is the Lorentz factor of the injected particles, and $n_{\pm}$ is the number density of positrons or electrons. Setting a minimum value for $\sigma_{*}$ yields a maximum value of the injected plasma density. With the parallel component of the velocity, we can control the injected current provided to the neutron star’s magnetosphere. Then, the magnetosphere extracts the sign of the charges required to drive the magnetospheric currents and sustain the global torsion of the magnetic field lines. The goal is to mimic the plasma supply provided by the vacuum work function and the pair-producing cascade close to the stellar surface, i.e. the polar cap gap, while still populating both the closed field line region and the return current. Previous works using global particle-in-cell simulations have already taken this approach as it is a computationally efficient way to populate the magnetosphere without requiring expensive quantum electrodynamics (QED) processes (Cerutti et al., 2015; Philippov et al., 2015b). Nevertheless, these processes are already implemented in the OSIRIS framework, to model QED cascades in future ultra-intense laser experiments (Grismayer et al., 2016, 2017) or in local neutron star polar caps (Cruz et al., 2021).
The second approach is to sustain a cold corotating atmosphere close to the surface, similarly to Hu and Beloborodov (2022); Bransgrove et al. (2022). We characterize the atmosphere by its value at the surface, $n_{0}$, and the scale height of the exponentially decaying profile, $\sigma_{atm}$:
@@
We ensure the non-depletion of this atmosphere up to $k_{atm}$ standard deviations from its base by injecting $n_{\mathrm{inj}}$ plasma number density every timestep when required. In the current study, we consider a plasma composed of electrons and positrons, the injection of ions is deferred to future studies. This approach differs from the previous one as it allows the magnetosphere to regulate the flux of plasma extracted from the corotating atmosphere. This extraction mechanism would then seed the cascading QED process. To reduce the computational cost of each simulation, we model the cascade with a heuristic pair production (Philippov et al., 2015a, b; Philippov and Spitkovsky, 2018; Chen et al., 2020; Guépin et al., 2020; Cruz et al., 2022, 2023). This model for the operation of the vacuum gap defines a threshold energy $\gamma_{thr}$ for a particle to pair produce. Particles that satisfy this condition can pair produce if their radial position corresponds to the range $R_{*}\leq r\leq R_{\mathrm{PP}}$. Also, this threshold energy has a polar dependence to inhibit pair production close to the polar axis, where the magnetic conversion does not occur due to the infinite magnetic field line curvature. The secondary pair is emitted at the same place and with the same momentum direction as the parent with a Lorentz factor of $\gamma_{pair}/2$. These parameters are user-defined and can be adjusted to study different magnetospheric solutions. This model has shown to be a good model for the polar cap and the outer gaps (Philippov et al., 2015a, b; Philippov and Spitkovsky, 2018; Chen et al., 2020; Guépin et al., 2020; Cruz et al., 2022, 2023).
@@ 2.8 Simulation initialization and transient phase
Simulations start in vacuum with a prescribed general-relativistic dipolar magnetic field (Rezzolla et al., 2001; Torres et al., 2023) threading the neutron star surface. Equivalently, the angular velocity of the star can be either in full rotation or gradually spun to its desired value. For the first scenario, the general-relativistic electric field (Torres et al., 2023) must be prescribed at $t=0$. In the second scenario, the rotation imparted by the stellar boundary condition (see Sect. 2.6.1) generates the self-consistent electric field by launching Alfvénic torsion waves (Cerutti et al., 2015). At $t=0$, we initiate the plasma injection mechanism described in Sect. 2.7. Due to the high magnetization, the extracted charges from the surface/atmosphere are accelerated along field lines, generating a poloidal current that opens up the field lines that cross the light-cylinder radius, located at $R_{LC}\equiv c/\Omega_{*}$. Beyond this distance, particles can no longer co-rotate with the field lines anchored to the stellar surface as they would need to co-rotate at speeds faster than the speed of light. These particles then sustain the torsion of the field lines that acquire a toroidal component. In the aligned rotator case, when the magnetic and rotation axes are aligned, the extracted current at the poles is negative due to the predominant outflow of electrons. The magnetosphere self-organizes to close the electric current, which generates the return current. This positive electric current neutralizes the charge density of the neutron star and flows between the closed and open field lines in a Y-shaped structure. Beyond the Y-point, the current sheet sustains the toroidal and poloidal magnetic field reversal at the equator. The magnetospheric solution starts to form in the back of the outwardly propagating Alfvénic torsion wave. We have assumed that after two rotation periods of the star, the magnetospheric solution has converged to the quasi-steady state.